### v1

@@ 6.2 Arbitrary Signal Space
We defer the proof to Section A.4. Our results imply that achieving SD-truthfulness using the EA mechanism is infeasible for arbitrary information structures beyond the binary-signal setting. Investigating alternative approaches, such as mechanism designs that leverage additional structure in the information space or stronger assumptions on agent strategies, remains an interesting direction for future work.
Although the EA mechanism is not SD-truthful when $|\Sigma|>2$, we show that it is still possible to obtain truthfulness as long as the information structure is known. We defer this discussion to Section A.5.
Remark 2 (Connection to Peer Truth Serum (Faltings et al., 2017)).
@@ 6.3 The Optimal Enforcement
We defer the proofs of Lemma 6.3 and 6.4 to Section A.6. While we present the analytical solution for the optimal enforcement to complete the theoretical framework, we note that this solution may be overfitted to the simplified effort model considered in our analysis. In practice, as demonstrated by our empirical results in the next section, enforcing the signal prior provides a strong approximation to the optimal enforcement and is our recommended approach.
7 Empirical Evaluations of SD-Truthful Mechanisms
In this section, we empirically compare the sensitivity of the proposed mechanisms across various settings. The results support our theoretical findings: EA consistently exhibits the highest sensitivity among SD-truthful mechanisms in the binary-signal setting (except when the signal prior is nearly uniform). We further evaluate each mechanism’s budgetary efficiency in an effort-elicitation scenario — measuring the minimum payment required to incentivize a target effort level. Again, EA proves its superiority, which has the best performance against any other SD-truthful and non-SD-truthful mechanisms. This supports our sensitivity analysis, demonstrating that a mechanism with high sensitivity can indeed lead to improved practical performance.
7.1 Datasets and Experiment Setup
We use two real-world datasets to estimate the information structure between two agents. The first dataset contains binary labels classifying whether a compound is appropriate or inappropriate to be synthesized (Baba et al., 2018). The second dataset collects the annotations of the sentiment of $300$ tweets, where the size of the signal space is 4 (Venanzi et al., 2015). We denote these two information structures as $J_{1}$ and $J_{2}$ respectively, and defer the details to Section C.2.
We consider the following mechanisms and their variants:
  • OA corresponds to the Output Agreement mechanism introduced in Section 4.1, which is truthful and SD-truthful only when signals are self-dominating.
@@ 7.1 Datasets and Experiment Setup
    • PTS-round represents the partition rounding reduction of PTS, which is SD-truthful. In particular, we normalize every individual score computed using reports from each question and average all individual scores.
@@
    • CA-partition represents the implementation where we get $K=n/2$ partitions of questions and score agents the average of the individual CA score for each partition (without normalization). By 4.3, CA-partition is SD-truthful when signals are binary.
    • CA-partition-round represents the implementation in which we not only partition the questions but also round each individual CA score. The final score is the average of the $K$ binary lottery. This implementation is SD-truthful for any size of $\Sigma$.
@@
  • MA corresponds to the original implementation of the Matching Agreement mechanism described in Appendix B, which is truthful but not SD-truthful.
    • MA-partition-round represents the partition rounding reduction of MA, which is SD-truthful.
  • We consider three implementations of EA.
    • EA-prior represents the Enforced Agreement mechanism with $\Phi=n\cdot\Pr(X_{a})$.
    • EA-uniform represents the Enforced Agreement mechanism with $\Phi=n\cdot[0.5,0.5]$.
    • EA-optimal — EA with the sensitivity-maximizing $\Phi$ computed according to Section 6.3.
For information structures with binary-signal settings, sensitivities of some of the mechanisms can be computed analytically as shown in Section 5.2.1. For general cases, we estimate sensitivity using a Monte Carlo approach. In particular, we simulate reports to each of the $n$ questions with both agents exerting full effort $e=1$ and run $\mathcal{M}$ to compute the scores. Repeating this process for $T$ times yields $T$ i.i.d. samples of $S^{\mathcal{M}}(e)$. We then compute the score of Alice when she deviates to a lower effort level at $e-\Delta e=0.8$, and obtain $T$ samples of $S^{\mathcal{M}}(e-\Delta e)$. The sensitivity of at effort $e$ is then estimated as $\frac{\mathbb{E}[S^{\mathcal{M}}(e)]-\mathbb{E}[S^{\mathcal{M}}(e-\Delta e)]}{\Delta e}\cdot\frac{1}{\operatorname{std}(S^{\mathcal{M}}(e))}$. This implies that we fix $e=1$ when estimating sensitivity in our experiments.
7.2 The Sensitivity of Discussed Mechanisms
We now compare the sensitivity of the discussed SD-truthful mechanisms. In Figure 1, we show the sensitivity when the prior of the binary signal is varied, i.e. $\Pr[X_{a}=0]\in[0.1,0.9]$, and the number of questions $n=100$. The details of how the prior is varied are provided in Section C.1. We observe that enforcing the signal prior in EA closely approximates the optimal enforcement. Remarkably, even EA-uniform, a detail-free version that doesn’t require knowledge of the prior, consistently outperforms any other SD-truthful mechanisms such as PTS-partition-round, which rely on prior knowledge for SD-truthfulness. While the partition rounding of PTS achieves sensitivity comparable to OA under a balanced prior (because PTS reduces to OA when $\Pr[X_{a}=0]$ is close to 0.5), its sensitivity drops significantly as the prior becomes more biased — precisely where EA demonstrates the highest sensitivity. Moreover, CA and MA exhibit identical sensitivities. In Appendix B, we formally prove that this pattern holds as long as signals are binary.
Figure 1: The sensitivity of SD-truthful mechanisms under various prior $\Pr[X_{a}=0]$.
Figure 2 present sensitivity as the number of questions $n\in\{10,20,\ldots,100\}$ increases, under information structures $J_{1}$ and $J_{2}$. As observed, the sensitivity of the mechanisms increases approximately proportional to $\sqrt{n}$, consistent with the theoretical prediction given in Eq. 2. Note that in $J_{2}$, where $|\Sigma|>2$, EA is no longer SD-truthful; nonetheless, we include its sensitivity for comparison.
(a) under $J_{1}$, $|\Sigma|=2$
(b) under $J_{2}$, $|\Sigma|=4$
Figure 2: A Comparison of the Sensitivity of Different Mechanisms. SD-truthful mechanisms are highlighted in bold for each setting, except for OA, which is SD-truthful only under self-dominating signals.
We additionally make the following observations. First, although EA is not SD-truthful in $J_{2}$ (but still truthful), its sensitivity is competitive with other truthful mechanisms. Second, the sensitivity of the partition rounding versions of CA and MA is consistently dominated by other SD-truthful mechanisms. This is largely due to their reliance on penalty questions, which do not contribute to the sensitivity, and their need for two questions to compute an individual score. Additionally, comparing the original implementation of CA and MA with their partition rounding variants, we observe that partition rounding can significantly degrade sensitivity. In particular, the sensitivity of CA-partition-round is just a quarter of the original. These findings motivate further exploration of more sensitive SD-truthful mechanisms beyond the binary-signal setting.
7.3 Budgetary Efficiency
We further evaluate the performance of the proposed mechanisms in an effort-elicitation setting, aiming to demonstrate that SD-truthful mechanisms with higher sensitivity lead to more efficient effort elicitation. This setting differs from the sensitivity estimation setup in several key ways. First, it involves more than two agents, with each pair potentially answering different subsets of questions. Second, we model a symmetric effort equilibrium, in contrast to the previous setting where Bob was assumed to always exert full effort. Finally, we incorporate effort costs and explicit payment schemes to better reflect realistic crowdsourcing conditions. These enhancements are intended to strengthen the robustness of our findings.
7.3.1 Model
Following a prior work (Zhang and Schoenebeck, 2023a), we consider an effort-elicitation model where each agent has an effort level $e_{i}$ with a convex cost function $c(e_{i})=e_{i}^{2}$ reflecting the increasing marginal cost of effort. The principal pays each agent a function of her score computed by a peer prediction mechanism, and agents best respond by choosing an effort that maximizes their expected utility — the difference between the expected payment and the cost of effort. Note that because the peer prediction score of an agent depends on all agents’ effort levels, we consider symmetric equilibrium as our solution concept, where every agent exerts the same effort level and no one wants to unilaterally deviate. This motivates the use of budgetary efficiency to measure the performance of a peer prediction mechanism, defined as the minimum cost of the budget to incentivize a desired effort at the symmetric equilibrium.
  • The linear payment scheme pays agent $i$ with $\pi_{i}=a\cdot S^{\mathcal{M}}_{i}+b$.
  • The winner-take-all tournament pays agent $i$ with $\pi_{i}=C_{\operatorname{WTA}}\cdot\mathbbm{1}[S^{\mathcal{M}}_{i}>S^{\mathcal{M}}_{j},~{}\forall j\neq i]$.
  • The Borda-count tournament pays agent $i$ with the number of agents she beats:
    $\pi_{i}=C_{\operatorname{Borda}}\cdot\#\text{beaten}=C_{\operatorname{Borda}}\sum_{j\neq i}\mathbbm{1}[S^{\mathcal{M}}_{i}>S^{\mathcal{M}}_{j}].$
In the above payment schemes, $a,b,C_{\operatorname{WTA}},C_{\operatorname{Borda}}$ are constant parameters that can be optimized for budgetary efficiency in practice given the score distributions. We further require payments to satisfy limited liability (i.e. the ex-post payment must be non-negative) and individual rationality (IR, i.e. the expected utility must be non-negative).
Note that both the peer prediction mechanism and the associated payment scheme influence budgetary efficiency. To preserve truthfulness in payments, a truthful mechanism must be paired with a linear payment scheme. In contrast, SD-truthful mechanisms can be combined with tournament-based payment schemes, which have been shown to be more efficient (Zhang and Schoenebeck, 2023a). Our goal is to investigate whether the combination of SD-truthful mechanisms and tournament payments can achieve greater efficiency than truthful mechanisms with linear payments.
(a) WTA Tournament, $e=0.7$
(b) WTA Tournament, $e=0.9$
(c) Borda-count Tournament, $e=0.7$
(d) Borda-count Tournament, $e=0.9$
Figure 3: Estimated Sensitivity v.s. Total Payment.
7.3.2 Experiment setup
We consider the information structure $J_{1}$ with 50 agents and 500 tasks, where each agent works on 50 tasks and each task is randomly assigned to 5 agents. To incentivize a desired effort level $e$ as a local symmetric equilibrium, we learn the parameters in the payment schemes by solving $\nabla\pi_{i}=\nabla c(e_{i})$ for any $i$. Note that $\nabla c(e_{i})$ can be directly computed given a desired effort $e_{i}=e$. To compute $\nabla\pi_{i}$, we repeat the same procedure as discussed in Section 7.1. In particular, we simulate samples of signals and apply various mechanisms to compute the peer prediction score $S_{i}^{\mathcal{M}}$ for each agent. Then, varying the effort from $e$ to $e-\Delta e$, we can estimate $\nabla S^{M}_{i}$ using a sample size $T=10,\!000$. Putting this estimation into the formula of various payment schemes thus gives us an estimation of $\nabla\pi_{i}$. After learning the parameters of the payment schemes, we compute the budgetary efficiency as the total payment to incentivize such an effort level.
7.3.3 Results
We compute the optimal linear payments — subject to limited liability and individual rationality — for all truthful mechanisms introduced in Section 7.1, and compare them against the tournament payments used with SD-truthful mechanisms. A lower total payment indicates greater budgetary efficiency. Figure 3 presents the results for SD-truthful mechanisms combined with winner-take-all and Borda-count tournaments. The red dashed line marks the lowest total payment required by any truthful mechanism under a linear payment scheme, while the blue dashed line indicates the minimum payment required to satisfy individual rationality. Each dot represents the total payment to elicit the desired effort for a given mechanism, ignoring the IR constraint.
We make several key observations. First, sensitivity correlates strongly with budgetary efficiency, consistent with the theoretical findings of prior work (Xu et al., 2024). Second, as predicted by our theoretical insights, the EA mechanism — particularly when enforcing the signal prior — consistently achieves the best performance. Third, the advantage of pairing an SD-truthful mechanism with a tournament payment scheme over a truthful mechanism with a linear payment scheme becomes more pronounced at higher target effort levels. This is because greater effort leads to more accurate peer prediction scores, thereby increasing sensitivity, which tournaments can leverage more effectively. These results highlight the practical importance of designing highly sensitive SD-truthful mechanisms, as they offer substantial improvements in the performance of incentive systems across real-world settings.
8 Conclusion and Future Work
This paper initiates the discussions of stochastically dominant-truthful peer prediction mechanisms, which ensure that in equilibrium, the score distribution under truth-telling first-order stochastically dominates any other strategy. Unlike traditional peer prediction mechanisms, which rely on linear utility assumptions, SD-truthful mechanisms can preserve truthfulness for any increasing utility function. We show that existing mechanisms fail to naturally achieve SD-truthfulness and propose a rounding method to enforce it. However, this process often reduces the sensitivity of the original mechanism, making it less efficient in practice. Additionally, we introduce the Enforcement Agreement (EA) mechanism, which is empirically shown to have the highest sensitivity among all SD-truthful mechanisms in the binary-signal setting.
Future work is needed to investigate more sensitive SD-truthful mechanisms beyond the binary-signal setting. A promising next step is to explore approximate SD-truthfulness, which allows truth-telling to lose a small utility for some utility functions. This relaxation could lead to mechanisms with better sensitivity and budgetary efficiency. Furthermore, one may generalize the concept of stochastically dominant strategy-proofness to other similar applications, such as prediction markets, proper scoring rules, and forecast competitions.
Appendix A Additional Proofs
A.1 The CA Mechanism
@@
Next, we reason about the distribution of scores under different strategies. Note that a single draw of the CA score can be $-1$, 00, or $1$. Therefore, to prove SD-truthfulness for a single draw of the CA score, it is sufficient to show that $\Pr\left(S^{CA}_{i}(\tau)=1\right)\geq\Pr\left(S^{CA}_{i}(\theta)=1\right)$ and $\Pr\left(S^{CA}_{i}(\tau)=-1\right)\leq\Pr\left(S^{CA}_{i}(\theta)=-1\right)$ for any strategy $\theta$. When signals are binary, there are only three types of untruthful strategy: 1) always reporting 0, 2) always reporting 1, and 3) always flipping 0 to 1 and 1 to 0. Any other strategy is a mixture of truth-telling and these strategies. Now, we show that all of these strategies are stochastically dominated by truth-telling by discussing two cases.
@@
A.2 The Partition Rounding Reduction
@@
  • Inductive step: suppose that for $K-1$ there exists some threshold $t_{K-1}$ such that
@@
The completes the proof of the necessity. ∎
A.3 SD-truthfulness for the Binary-signal Setting
@@
To prove SDT, we have to show that truth-telling results in a score distribution that first-order stochastic dominates the score distribution under any other strategy. Suppose W.L.O.G. that $m_{0}>n_{0}$ and the mechanism will randomly flip Alice’s reports on $m_{0}-n_{0}$ questions from 0 to 1. If Alice reports truthfully, there are $n_{0}$ questions where Alice reports 00 and will be scored 1 if Bob also reports 00; there are $m_{1}$ questions where Alice reports $1$ and will be scored 1 if Bob also reports $1$; and there are $m_{0}-n_{0}$ questions where Alice observes 00 but her reports are flipped to 1 which means that she will be scored 1 if Bob reports $1$ on those questions. Therefore, the final score of Alice can be expressed as
$\ n\cdot S(\tau)\sim{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathrm{Bin}(n_{0},p_{00})}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathrm{Bin}(m_{1},p_{11})}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathrm{Bin}(m_{0}-n_{0},p_{01})}.$ (5)
@@
$n\cdot S(\theta){\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\sim\mathrm{Bin}(n_{0}-k,p_{00})}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathrm{Bin}(m_{1}-k,p_{11})}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathrm{Bin}(m_{0}-n_{0}+k,p_{01})}+{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\mathrm{Bin}(k,p_{10})}.$ (6)
Compared with the truth-telling score distribution in Eq. 3, any untruthful reporting can be viewed as subtracting $k$ samples from ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathrm{Bin}(n_{0},p_{00})}$ and draw $k$ more samples from ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\mathrm{Bin}(k,p_{10})}$; subtracting $k$ samples from ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathrm{Bin}(m_{1},p_{11})}$ and draw $k$ more samples from ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathrm{Bin}(m_{0}-n_{0}+k,p_{01})}$.
@@
A.4 SD-truthfulness for the Three-signal Setting
Proof of 6.2.

### v2

@@ C.1 Arbitrary Signal Space
We defer the proof to Section F.5. Our results imply that achieving SD-truthfulness using the EA mechanism is infeasible for arbitrary information structures beyond the binary-signal setting. Investigating alternative approaches, such as mechanism designs that leverage additional structure in the information space or stronger assumptions on agent strategies, remains an interesting direction for future work.
C.2 Truthfulness of EA
Although the EA mechanism is not SD-truthful when $|\Sigma|>2$, we show that it is still possible to obtain truthfulness as long as the information structure is known.
$\max_{\vartheta}\quad$ $\sum_{i,j\in\Sigma}\vartheta_{i,j}\;p_{ij}$
$\vartheta_{i,j}\in\mathbb{Z}_{\geq 0},\quad\forall i,j\in\Sigma.$
Proposition C.3.
Remark C.4 (Connection to Peer Truth Serum Faltings et al. [2017]).
@@ C.3 The Optimal Enforcement
We defer the proofs of Lemma C.5 and C.6 to Section F.6.
Appendix D The Matching Agreement Mechanism
The matching agreement mechanism is a follow-up of the correlated agreement mechanism aiming to address more complex cheating strategies Zhang and Schoenebeck [2023b]. Similar to CA, which determines “agreement” using the delta matrix (see Section 4.3), MA determines “agreement” using a different method depending on the information structure.
Definition D.1.
Let $\Gamma$ be a $|\Sigma|\times|\Sigma|\times|\Sigma|$ tensor where
Next, the mechanism randomly samples a bonus question $j$ and a penalty question $k$ uniformly at random. An individual score computed using these two questions is thus $S^{MA}_{i}=T_{\Gamma}(\hat{X}_{a,j},\hat{X}_{b,j},\hat{X}_{b,k})$. In words, the matching agreement mechanism awards Alice a score of 1 if, when predicting her response on question $b$, the posterior conditioned on Bob’s response to the same question provides a better prediction than the posterior conditioned on Bob’s response on a different question $q$. Similar to CA, each bonus question can be paired with n1n-1 penalty questions. Therefore, the final score of MA is the average of $n\cdot(n-1)$ individual scores.
The Sensitivity of MA Under the Partition Rounding Reduction
We further investigate the sensitivity of the MA mechanism under the partition reduction. The MA mechanism requires two questions to compute an individual score, implying that $K=n/2$. Clearly, $S_{\inf}=0$ and $S_{\sup}=1$ by definition. By Equation 1, to compute the sensitivity, we only need to compute the expected individual score $m_{i}(e)$ and its derivative. When Alice exerts full effort, the probability of obtaining a score of 1 can be computed as follows:
where $(x)^{+}=x$ if $x>0$ and 0 otherwise.
When Alice exerts no effort,
$=\frac{1}{2}$
When signals are binary, we find that MA reduces to CA.
Lemma D.2.
Proof.
$=\sum_{\sigma_{1},\sigma_{2}}(\Delta_{\sigma_{1},\sigma_{2}})^{+}$
An immediate consequence of Lemma D.2 is that the expected score under MA is a linear transformation of the expected score under CA. This further implies that the partition rounding reductions of MA and CA has the same sensitivity under the binary-signal setting.
Proposition D.3.
Proof.
By Lemma D.2, we know that
$\nabla m_{i}^{MA}(e)=\frac{1}{2}\nabla m_{i}^{CA}(e).$
By Equation 1,
$=\delta^{CA}(e).$
Appendix E Experiments (Continue)
All codes for our experiments are provided in https://github.com/DavidXu999/Stochastically-Dominant-Peer-Prediction.
E.1 Information Structure
For Section 7.2, we consider two real-world crowdsourcing datasets, each having a signal space of size $|\Sigma_{1}|=2$ and $|\Sigma_{2}|=4$. The information structure between a pair of agents can be characterized by a Dawid-Skene model Dawid and Skene [1979]. In particular, each task has an underlying ground truth $\omega$ that is i.i.d. sampled from a prior distribution $W$. Conditioning on $\omega$, each agent receives a signal $\sigma$ according to conditioned distribution $\Gamma$, where $\Gamma_{i,j}=\Pr[\sigma=j\mid\omega=i]$.
In the first dataset [Baba et al., 2018], $18$ agents each labels whether a compound is appropriate or inappropriate to be synthesized. The prior of the ground truth and the conditioned signal distribution are
In the second dataset Venanzi et al. [2015], $110$ agents provide the annotations of the sentiment of $300$ tweets. There are four ground truth labels and four possible signals where the information structure is
E.2 The Sensitivity of Original Implementations
How much does the partition rounding reduction harm the sensitivity? We compare the partition rounding reduction implementations of CA, MA, and PTS with their original implementations in Figure 3. To be clear, other than the implementations considered in Section 7, we consider the following implementation of the mechanisms.
@@ E.2 The Sensitivity of Original Implementations
  • CA-partition represents the implementation where we get $K=n/2$ partitions of questions and score agents the average of the individual CA score for each partition (without normalization). By F.1, CA-partition is SD-truthful when signals are binary.
@@
  • MA corresponds to the original implementation of the Matching Agreement mechanism described in Appendix D, which is truthful but not SD-truthful.
(a) Mechanisms shown in Section 7 under $J_{1}$
(b) Other Mechanisms under $J_{1}$
(c) Mechanisms shown in Section 7 under $J_{2}$
(d) Other Mechanisms under $J_{2}$
Figure 3: A Comparison of the Sensitivity of Different Mechanisms.
We observe that, relative to the original implementations (solid lines), the partition-rounding reductions (dashed lines) consistently exhibit lower sensitivity. This observation further demonstrates the superior sensitivity of EA mechanisms, as it has the property of SD-truthfulness in the binary setting with no need for partition rounding reduction.
Appendix F Additional Proofs
F.1 The CA Mechanism
@@
Next, we reason about the distribution of scores under different strategies. Note that a single draw of the CA score can be 1-1, 0, or $1$. Therefore, to prove SD-truthfulness for a single draw of the CA score, it is sufficient to show that $\Pr\left(S^{CA}_{i}(\tau)=1\right)\geq\Pr\left(S^{CA}_{i}(\theta)=1\right)$ and $\Pr\left(S^{CA}_{i}(\tau)=-1\right)\leq\Pr\left(S^{CA}_{i}(\theta)=-1\right)$ for any strategy $\theta$. When signals are binary, there are only three types of untruthful strategy: 1) always reporting 0, 2) always reporting 1, and 3) always flipping 0 to 1 and 1 to 0. Any other strategy is a mixture of truth-telling and these strategies. Now, we show that all of these strategies are stochastically dominated by truth-telling by discussing two cases.
@@
F.2 The Direct Rounding Reduction
Proof of 5.2.
$\Rightarrow\quad$ $\tilde{\delta}^{\mathcal{M}}(e)=\frac{\nabla m^{\mathcal{M}}(e)}{\sqrt{(m^{\mathcal{M}}(e)-S^{\mathcal{M}}_{\inf})\left(S^{\mathcal{M}}_{\sup}-m(e)\right)}}.$ (3)
F.3 The Partition Rounding Reduction
@@
  • Inductive step: suppose that for K1K-1 there exists some threshold $t_{K-1}$ such that
@@
This completes the proof of the necessity. ∎
F.4 EA in the Binary-signal Setting
@@
To prove SDT, we have to show that truth-telling results in a score distribution that first-order stochastic dominates the score distribution under any other strategy. Suppose W.L.O.G. that $m_{0}>n_{0}$ and the mechanism will randomly flip Alice’s reports on $m_{0}-n_{0}$ questions from 0 to 1. If Alice reports truthfully, there are $n_{0}$ questions where Alice reports 0 and will be scored 1 if Bob also reports 0; there are $m_{1}$ questions where Alice reports $1$ and will be scored 1 if Bob also reports $1$; and there are $m_{0}-n_{0}$ questions where Alice observes 0 but her reports are flipped to 1 which means that she will be scored 1 if Bob reports $1$ on those questions. Therefore, the final score of Alice can be expressed as
$\ n\cdot S(\tau)\sim{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathrm{Bin}(n_{0},p_{00})}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathrm{Bin}(m_{1},p_{11})}+{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\mathrm{Bin}(m_{0}-n_{0},p_{01})}.$ (4)
@@
$n\cdot S(\theta){\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\sim\mathrm{Bin}(n_{0}-k,p_{00})}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathrm{Bin}(m_{1}-k,p_{11})}+{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\mathrm{Bin}(m_{0}-n_{0}+k,p_{01})}+{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\mathrm{Bin}(k,p_{10})}.$ (5)
Compared with the truth-telling score distribution in Equation 2, any untruthful reporting can be viewed as subtracting $k$ samples from $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathrm{Bin}(n_{0},p_{00})$} and draw $k$ more samples from $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\mathrm{Bin}(k,p_{10})$}; subtracting $k$ samples from $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathrm{Bin}(m_{1},p_{11})$} and draw $k$ more samples from $\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\mathrm{Bin}(m_{0}-n_{0}+k,p_{01})$}.
@@
F.5 EA in the Three-signal Setting
Proof of C.2.